OpenStudy (anonymous):

when you multiply two numbers with like signs the answer is always blank give two examples

TheSmartOne (thesmartone):

the answer is always positive -3x-3=+9 5x5=+25

TheSmartOne (thesmartone):

@mcnabkt

OpenStudy (anonymous):

ohhh okay well thanks

TheSmartOne (thesmartone):

your welcome.

OpenStudy (anonymous):

When you multiply two numbers with unlike signs the answer is always

OpenStudy (anonymous):

oh would this be negative?

OpenStudy (anonymous):

yup it would be negative

OpenStudy (anonymous):

two examples?

OpenStudy (anonymous):

-8 x3= -21 12 x -4= -48

TheSmartOne (thesmartone):

5x-3=-15 6x-4=-24

OpenStudy (anonymous):

i need another positive one

OpenStudy (anonymous):

nvm